OpenStudy (anonymous):

Of the 630 at South School, three fifths are girls. Four out of every 7 boys can swim . How many boys can swim?

OpenStudy (anonymous):

if three fifths are girls then two fifth are boys, so: 2/5 * 630 = number of boys = 252 boys 4/7 = ratio of boys that can swim, so: 4/7 * 252 = ratio of boys that can swim= 144 therefore 144 boys can swim! YaaaaaaaaaaaaaaY

OpenStudy (anonymous):